OpenStudy (anonymous):

Explain why f(x)=(sqare root x-5) / x is not a rational function

OpenStudy (anonymous):

ANY FUNCTION UNDER SQUARE ROOT CAN NEVER BE A RATIONAL FUNCTION SO IT IS A IRRATIONAL FUNCTION

OpenStudy (anonymous):

What if the equation is f(x)= 10x / (square root x - (52t + 3))

OpenStudy (anonymous):

since the suare root is on bottom, is that still not a rational equation?

OpenStudy (lgbasallote):

why can't a function under a square root be a rational function @rvgupta ?

OpenStudy (lgbasallote):

anyway...is this the question @stardust3829 \[\frac{\sqrt{x-5}}x\]or\[\frac{\sqrt x - 5}x\]

OpenStudy (anonymous):

the top one

OpenStudy (lgbasallote):

...and who said it wasn't rational?

OpenStudy (anonymous):

how do you do the square root on the computer? my teacher

OpenStudy (lgbasallote):

can you see a button below the reply box that says \(\boxed{\Sigma \;\text{Equation}}\). click it you'll see the square root

OpenStudy (lgbasallote):

anyway...do you know the definition of a rational function?

OpenStudy (anonymous):

the quotient of two polynomials

OpenStudy (lgbasallote):

right. and do you know the definition of a polynomial?

OpenStudy (anonymous):

two or more monomials

OpenStudy (lgbasallote):

not exactly...a monomial is still a polynomial (monomial, binomials, trinomials, etc are types of polynomials)

OpenStudy (anonymous):

never mind then

OpenStudy (lgbasallote):

a polynomial is an expression wherein ALL the exponents of every term are nonnegative integer exponents a square root has an exponent of 1/2..and 1/2 is not an integer..therefore, a square root is not a polynomial

OpenStudy (anonymous):

and because it isn't a polynomial, it cant be the quotient of two polynomial, thus not a rational function...right

OpenStudy (lgbasallote):

right